# Critical Metallicity of Cool Supergiant Formation. II. Physical Origin

### v1

@@ 3.3 Grid (d): Shell Hydrogen Burning Rate and Opacity
The reaction rates of H burning can significantly influence whether a star evolves into an RSG; however, it remains unclear whether this effect originates primarily from core H burning during the main-sequence phase or from shell H burning in the post-main-sequence phase. To address this question, we constructed another set of models, Grid (d), in which the CNO cycle reaction rates are modified only during the post-main-sequence shell H-burning phase using $\eta_{\rm shellCNO}$, while remaining unchanged during the main sequence.
Figure 3 shows the resulting $R_{\rm max}$ values for models with varying $\eta_{\rm shellCNO}$. Despite some scatter in the data, it is clear that $\eta_{\rm shellCNO}$ has little impact on the final radii of supergiants. In general, even when $\eta_{\rm shellCNO}$ is extremely low, stars with sufficiently high opacity can still expand to RSG dimensions.
This suggests that the strong dependence of $R_{\rm max}$ on $\eta_{\rm CNO}$ observed in Grid (c) primarily originates from core H burning during the main sequence, rather than from shell H burning in the post-main-sequence phase. In other words, our results reveal that nuclear burning during the main sequence can significantly affect the degree of envelope expansion during the post-main-sequence stage.
3.4 Grid (e): Helium Burning Rate and Opacity
@@
In Ou & Chen (2024), we showed that enhanced He burning can actually suppress the expansion toward the RSG phase. To further demonstrate this effect, we construct a dedicated set of models, Grid (e), in which the triple-alpha reaction rate is scaled by a linear coefficient $\eta_{3\alpha}$.
Figure 4 displays the Grid (e) with varying $\eta_{3\alpha}$ and $\zeta_{\rm opacity}$. In the lower-right region of the diagram, some models with low $\zeta_{\rm opacity}$ and high $\eta_{3\alpha}$ fail to evolve into RSGs. An increased $\eta_{3\alpha}$ not only fails to promote RSG formation but may even inhibit it, particularly at low opacity.
In summary, we conducted a series of experiments to investigate the effects of opacity and nuclear reaction rates on the formation of RSGs. Our results show that both metallicity-dependent factors, opacity and the H-burning rate during the main-sequence phase, play important roles in promoting the evolution into an RSG, whereas enhanced core He-burning tends to suppress it.
Figure 5: Effective temperature ($T_{\rm eff}$), central temperature ($T_{\rm c}$), surface pressure ($P_{\rm surf}$), and central pressure ($P_{\rm c}$) at both the ZAMS and TAMS for $25\,M_{\odot}$ models with $Z=0.001$ from Grid (c), where the CNO cycle energy generation rate is scaled by $\eta_{\rm CNO}$ and the opacity parameter by $\zeta_{\rm opacity}$. In each panel, models are plotted are plotted as functions of $\eta_{\rm CNO}$, and the color of each point represents the maximum stellar radius ($R_{\rm max}$) reached during the core He-burning phase, with red indicating RSG-scale radii and blue indicating BSG-scale radii.
@@
Figure 6: ZAMS radii (left panels) and TAMS radii (middle and right panels) for $25\,M_{\odot}$ models with $Z=0.001$ from Grid (c), systematically varying the CNO cycle energy generation rate (scaled by $\eta_{\rm CNO}$) and the opacity parameter ($\zeta_{\rm opacity}$). The top row shows models arranged along $\eta_{\rm CNO}$, and the bottom row along $\zeta_{\rm opacity}$. In each panel, colors indicate the maximum radius reached during the core He-burning phase ($R_{\rm max}$), with red corresponding to RSG sizes and blue to BSG sizes. The middle panels display individual models as scattered points at TAMS, while the right panels present smoothed contour plots using the same dataset. Notably, models with TAMS radii $\gtrsim 12\,R_{\odot}$ predominantly evolve into RSGs, whereas those below this threshold remain BSGs.
4 Direct Indicator for RSG Formation
The results in Section 3 indicate that both opacity and core H-burning rates contribute to promoting stellar evolution into the RSG phase. In this section, we aim to develop a unified picture that explains how these two factors jointly determine whether a star evolves into a red supergiant.
4.1 Threshold for Evolution into the RSG Phase at a Constant Mass
One key finding from the previous section is that the H-burning rate during the main-sequence phase significantly influences the star’s post-main-sequence evolution. To investigate this delayed effect of main-sequence processes on post-main-sequence evolution, we analyze how the stellar parameters in the main sequence correlate with the maximum radius ($R_{\rm max}$) reached during the core He-burning stage. In particular, we examine two key moments in the history of stellar evolution: the ZAMS and the terminal-age main sequence (TAMS).
To avoid the complication of varying stellar mass, we begin with Grid (c), in which the initial mass is fixed at $25\,M_{\odot}$. Figure 5 shows the effective temperature ($T_{\rm eff}$), central temperature ($T_{\rm c}$), surface pressure ($P_{\rm surf}$), and central pressure ($P_{\rm c}$) at both the ZAMS and TAMS for each model. The data points are plotted as functions of $\eta_{\rm CNO}$, and colored by the resulting $R_{\rm max}$. Our main goal is to identify the boundary separating models that evolve into RSGs (red points) from those that remain as BSGs (blue points), and to explore how this transition correlates with the physical parameters at ZAMS and TAMS. If a threshold value in one of these parameters can clearly distinguish the red and blue points, it may provide an important clue to the underlying evolutionary mechanism.
At ZAMS, while RSG and BSG models show a clear visual separation, the dividing line does not align with any fixed value of $T_{\rm eff}$, $T_{\rm c}$, $P_{\rm surf}$, or $P_{\rm c}$. Thus, these ZAMS parameters alone do not serve as direct predictors of whether a star will become an RSG.
[…]
$T_{\rm eff}$ is tightly related to the stellar radius as the envelope cools during expansion. Thus, we further examine how the radii at the ZAMS and TAMS correlate with the maximum radius, $R_{\rm max}$. In Figure 6, we present the ZAMS and TAMS radii as functions of $\eta_{\rm CNO}$ and $\zeta_{\rm opacity}$, with colors of data points indicating the corresponding values of $R_{\rm max}$.
At the ZAMS, when core H burning has just begun, the stellar radius ($R_{\rm ZAMS}$) shows minimal dependence on $\eta_{\rm CNO}$, but displays moderate variation with $\zeta_{\rm opacity}$. At this stage, the boundary separating RSG and BSG data points appears as a tilted line rather than aligning with any constant radius. This suggests that the ZAMS radius alone is not a robust predictor of whether a star will eventually evolve into an RSG.
As shown in the middle and right panels of Figure 6, by the TAMS, after the main-sequence evolution is complete, the stellar radius ($R_{\rm TAMS}$) exhibits a much stronger dependence on $\eta_{\rm CNO}$, while opacity also contributes to increasing the radius to some extent, particularly for $\zeta_{\rm opacity}\gtrsim 10^{-3}$. RSG and BSG models are now clearly divided by a horizontal boundary. Specifically, if the TAMS radius exceeds a threshold of $\sim 12\,R_{\odot}$, a $25\,M_{\odot}$ star eventually evolves into an RSG; otherwise, it remains a BSG. These results indicate that the stellar radius attained by the end of the main-sequence phase is a decisive factor for whether a star will ultimately evolve into an RSG.
To assess whether this threshold radius correlates with the He core radius, we plot the distribution of TAMS stellar radius versus the He core radius at its formation, with point colors indicating $R_{\rm max}$. We find that, for any given TAMS radius, the models span a broad range of core radii. The boundary separating RSGs (red points) from BSGs (blue points) at $R_{\rm TAMS}\sim 12\,R_{\odot}$ does not align with a unique He core radius, but spans a broad range of core sizes. This confirms that whether a star evolves into an RSG is not determined by He core radius, but rather by the radius of the entire star at the onset of the post-main-sequence phase.
Figure 7: Distribution of TAMS radii and He core radii at its formation ($R_{\rm c}$), with colors indicating the maximum radius during core He burning ($R_{\rm max}$). The boundary between RSG and BSG models aligns closely with a horizontal line of constant TAMS radius, rather than with the He core radius.
These findings point to a unified principle of the metallicity effect on supergiant evolution. At a fixed initial mass, a star can eventually evolve into an RSG if its radius at the onset of post-main-sequence stage exceeds a certain threshold. Conversely, if the star fails to reach this threshold, it will remain a BSG. Both opacity and the core H-burning rate are influenced by metallicity and affect the stellar structure during the main-sequence phase. Together, they shape the stellar radius at the TAMS, which ultimately determines whether the star will evolve into an RSG.
We note that altering the core He-burning rates during the post-main-sequence phase, as explored in Grid (e), can still influence the final evolutionary outcome. However, it does not account for the systematic trend in which higher metallicity favors RSG formation over BSGs. For this reason, we exclude this effect from our main analysis.
4.2 Threshold TAMS Radii for Different Masses
@@
To investigate how stellar mass influences the threshold radius required for RSG formation, we conduct a similar analysis using Grid (a). We examine the boundary between red and blue data points, which delineates the transition between stars that evolve into RSGs and those that remain BSGs. As shown in the left panel of Figure 8, this boundary is no longer located at a constant TAMS radius but increases systematically with stellar mass.
To obtain a coarse fit for the threshold TAMS radius required for RSG formation as a function of mass, we rescale the TAMS radius ($R_{\rm TAMS}$) by $M_{\rm TAMS}^{\alpha}$, where $M_{\rm TAMS}$ is the stellar mass at the TAMS and $\alpha$ is a trial exponent. By manually adjusting $\alpha$, we identify a scaling relation that yields an approximately horizontal boundary separating RSG and BSG cases. As shown in the right panel of Figure 8, this leads to the following approximate criterion for RSG formation:
@@
This relation shows the approximate mass dependence of the threshold radius for RSG formation. Note that $M_{\rm TAMS}$ differs only slightly from the initial mass, as mass loss during the main sequence is typically modest.
We propose the TAMS radius as a practical criterion for determining whether a star evolves into the RSG phase. However, the underlying physical mechanism responsible for this threshold, and its connection to metallicity, requires further investigation. In the following sections, we address two key questions:
(1) How does metallicity affect the stellar radius?
@@
We begin by quantifying how metallicity affect stellar structure and evolution in Section 5, followed by an classification of the evolutionary pathways of supergiants in Section 6. Based on these pathways, we then analyze how the TAMS radius regulates supergiant expansion in Section 7.
@@
5 Metallicity Effect on Stellar Radius
@@
To evaluate how metallicity affects stellar radius, we begin by analyzing evolutionary tracks from Grids (a) and (c), and then apply homology relations to perform a dimensional analysis.
5.1 Metallicity Dependence of Stellar Evolutionary Tracks
In previous sections, we demonstrated that both $\eta_{\rm CNO}$ and $\zeta_{\rm opacity}$ can influence whether a star evolves into an RSG by modifying its TAMS radius. To investigate how these parameters affect stellar evolution in detail, we examine evolutionary tracks on the HR diagram. We focus on two subsets of models from Grid (c): one in which $\eta_{\rm CNO}$ is varied while $\zeta_{\rm opacity}$ is held fixed, and the other in which $\zeta_{\rm opacity}$ is varied while $\eta_{\rm CNO}$ is fixed.
Figure 9 presents the HR diagram for models with an initial mass of $25\,M_{\odot}$ and metallicity $Z=0.001$, where $\zeta_{\rm opacity}$ is varied, and the reaction rate of the CNO cycle is fixed at its standard value ($\eta_{\rm CNO}=1$). Changes in $\zeta_{\rm opacity}$ lead to noticeable differences in luminosity starting from the ZAMS, along with moderate variations in $T_{\rm eff}$. In general, a higher opacity (larger $\zeta_{\rm opacity}$) results in a lower luminosity, which would suggest a smaller radius. However, higher opacity also reduces $T_{\rm eff}$, partially offsetting the luminosity-driven effect on radius. ZAMS positions lie nearly parallel to iso-radius lines, indicating weak radius sensitivity to opacity at this stage. As the stars evolve, the relationship between opacity and radius shifts. By the time they reach the TAMS, just before the expansion into the supergiant phase, models with higher opacity display systematically larger radii.
@@ 5.1 Metallicity Dependence of Stellar Evolutionary Tracks
Figure 10 displays the HR diagram for another subset of $25\,M_{\odot}$ models at $Z=0.001$, this time with fixed opacity ($\zeta_{\rm opacity}=0.001$, the standard value for this metallicity) and varying $\eta_{\rm CNO}$. In this case, the ZAMS positions are not significantly affected by changes in $\eta_{\rm CNO}$, but the subsequent evolutionary tracks exhibit a systematic shift in $T_{\rm eff}$ across different values of $\eta_{\rm CNO}$. Throughout the main-sequence phase and up to the TAMS, higher values of $\eta_{\rm CNO}$ result in lower $T_{\rm eff}$, while the luminosities remain similar. Consequently, stars with enhanced CNO-cycle rates attain larger radii by the TAMS.
These results show that during the main-sequence phase, both increased opacity and a higher CNO-cycle reaction rate shift the stellar evolutionary tracks toward cooler effective temperatures and larger radii on the HR diagram. At higher metallicities, where both the opacity and the CNO-cycle rate are elevated, stars are thus expected to develop systematically larger TAMS radii.
@@
To examine the cumulative effect of metallicity on TAMS radius, Figure 11 presents evolutionary tracks of $25\,M_{\odot}$ models with varying metallicities from Grid (a). While the ZAMS positions show a relatively irregular distribution across different $Z$, the subsequent evolutionary tracks exhibit a more systematic trend toward the TAMS. In general, higher-metallicity models tend to reach lower luminosities but also lower $T_{\rm eff}$ at the TAMS, resulting in systematically larger stellar radii. This trend reflects the combined influence of metallicity on stellar evolution through its effects on both opacity and nuclear burning.
Figure 9: Evolutionary tracks of $25\,M_{\odot}$, $Z=0.001$ stars with different $\zeta_{\rm opacity}$ on the HR diagram. Dashed lines indicate constant-radius contours at 10, 100, and $1000\,R_{\odot}$. Black triangles mark the ZAMS, black circles indicate the TAMS, and black stars denote the maximum stellar radius reached during the core He-burning phase. Tracks that evolve into RSGs are shown in red, while those that remain BSGs are shown in blue.
@@
Figure 10: Evolutionary tracks of $25\,M_{\odot}$, $Z=0.001$ stars with different $\eta_{\rm CNO}$ on the HR diagram. Dashed lines indicate constant-radius contours at 10, 100, and $1000\,R_{\odot}$. Black triangles mark the ZAMS, black circles indicate the TAMS, and black stars denote the maximum stellar radius reached during the core He-burning phase. Tracks that evolve into RSGs are shown in red, while those that remain BSGs are shown in blue.
@@
Figure 11: Evolutionary tracks of $25\,M_{\odot}$ stars with different metallicities on the HR diagram. Dashed lines indicate constant-radius contours at 10, 100, and $1000\,R_{\odot}$. Black triangles mark the ZAMS, black circles indicate the TAMS, and black stars denote the maximum stellar radius reached during the core He-burning phase. Tracks that evolve into RSGs are shown in red, while those that remain BSGs are shown in blue.
5.2 Homology Relation between Metallicity and Radius

### v2

@@ III.3 Grid (d): Shell Hydrogen Burning Rate and Opacity
The reaction rates of H burning can strongly influence whether a star evolves into an RSG, but it is unclear whether this sensitivity originates mainly from core or shell H burning. One might intuitively expect shell H burning in the post-main-sequence phase to drive envelope expansion. To test this, we construct Grid (d), where the CNO reaction rates are modified only during post-main-sequence shell H burning (scaled by $\eta_{\rm shellCNO}$) while leaving main-sequence burning unchanged.
Figure 3 presents the resulting $R_{\rm max}$ values. Despite some scatter, $\eta_{\rm shellCNO}$ has only a minor effect on the maximum radii of supergiants. This shows that the strong dependence of $R_{\rm max}$ on $\eta_{\rm CNO}$ seen in Grid (c) may not originate from shell H burning, but rather from modifications to core H burning during the main sequence. Thus, we obtain the intriguing result that nuclear burning in the main sequence can significantly influence the degree of envelope expansion later in evolution.
III.4 Grid (e): Helium Burning Rate and Opacity
@@
In Ou & Chen (2024), we showed that enhanced He burning can actually suppress the expansion toward the RSG phase. To further demonstrate this effect, we construct Grid (e), in which the triple-alpha reaction rate is scaled by a linear coefficient $\eta_{3\alpha}$, as shown in Figure 4. In the lower-right region of the diagram, some models with low $\zeta_{\kappa}$ and high $\eta_{3\alpha}$ fail to evolve into RSGs. An increase in $\eta_{3\alpha}$ not only does not promote RSG formation but can even inhibit it, particularly at low opacity.
In summary, we conducted a series of experiments and find that both higher opacity and a stronger core H-burning rate during the main sequence favor expansion into the RSG regime, whereas enhanced core He burning suppresses it. These trends explain why higher-metallicity stars more readily become RSGs. However, although opacity and core H burning jointly shape the outcome, the common underlying trigger that enables the RSG transition remains unclear, which motivates further investigation in the next section.
@@
Figure 5: Effective temperature ($T_{\rm eff}$), central temperature ($T_{\rm c}$), surface pressure ($P_{\rm surf}$), central pressure ($P_{\rm c}$), and stellar radius ($R_{*}$) at the ZAMS and TAMS for $25\,M_{\odot}$ models from Grid (c). In each panel, the ZAMS value is shown on the X-axis and the TAMS value on the Y-axis. Points are colored by the maximum stellar radius ($R_{\rm max}$) reached during core He burning, with red indicating RSG-scale radii and blue indicating BSG-scale radii. The bottom-right panel shows a smoothed contour representation of the same $R_{*}$ data displayed in the bottom-left panel. Notably, models with TAMS radii $\gtrsim 12\,R_{\odot}$ predominantly evolve into RSGs, whereas those below this threshold remain BSGs.
IV Threshold for RSG Formation
We now aim to develop a unified picture of how opacity and core H-burning rates jointly modify the stellar structure and therefore determine whether a star evolves into an RSG. Specifically, we seek to identify the key structural condition established during the main-sequence phase that ultimately enables the formation of RSG.
To this end, we examine stellar parameters at two evolutionary points—the ZAMS and the terminal-age main sequence (TAMS)—and assess how they correlate with the maximum radius ($R_{\rm max}$) reached during core He burning. In our MESA models, the ZAMS is defined using the "create_zams" template, which generates a ZAMS model that serves as the initial condition for subsequent evolution. Accordingly, we extract the ZAMS parameters from the first timestep of each evolutionary track. The TAMS is defined as the last timestep before the emergence of a He core, identified by the onset of a non-zero He-core mass in the MESA output.
Figure 5 uses Grid (c), in which the initial mass is fixed at $25\,M_{\odot}$, to show how ZAMS and TAMS stellar properties correlate with the supergiant outcome. Each panel displays one stellar parameter—effective temperature ($T_{\rm eff}$), central temperature ($T_{\rm c}$), surface pressure ($P_{\rm surf}$), central pressure ($P_{\rm c}$), or stellar radius ($R_{*}$)—with its ZAMS value on the X-axis and its TAMS value on the Y-axis. Each model in Grid (c) is plotted on these panels and colored by its resulting $R_{\rm max}$.
Our aim is to identify the boundary separating models that evolve into RSGs (red points) from those that remain BSGs (blue points). The results show that the red–blue division is well defined in the surface parameters ($T_{\rm eff}$, $P_{\rm surf}$, and $R_{*}$), but not in the core parameters ($T_{\rm c}$ and $P_{\rm c}$). Moreover, these division lines appear nearly horizontal, aligning with fixed TAMS surface values rather than ZAMS values. In particular, the TAMS radius emerges as a strong indicator of RSG formation: stars with any ZAMS radius can become RSGs, but only those whose TAMS radius exceeds a threshold of $\sim 12\,R_{\odot}$ ultimately evolve into RSGs.
@@ IV Threshold for RSG Formation
Figure 6 uses Grid (a) to conduct a similar analysis for stars with different initial masses. The red-blue boundary is no longer located at a constant TAMS radius but increases systematically with stellar mass. As shown in the right panel, we obtain an approximate criterion for RSG formation:
@@
where $R_{\rm TAMS}$ and $M_{\rm TAMS}$ are the radius and the mass, respectively, at the TAMS. Note that $M_{\rm TAMS}$ differs only slightly from the initial mass ($M_{\rm i}$), as mass loss during the main sequence is typically modest.
Therefore, we arrive at a unified picture of how opacity and the core H-burning rate promote RSG formation: both factors modify the stellar structure during the main-sequence phase, leading to different TAMS radii, which in turn serve as the decisive quantity for RSG formation. Regardless of which process alters the TAMS radius, a star can evolve into an RSG as long as its TAMS radius exceeds a critical threshold. However, the physical mechanism linking metallicity, the TAMS radius, and RSG formation remains unclear. In the following sections, we address two key questions:
(1) How does metallicity affect the TAMS radius?
@@
We begin by quantifying how metallicity affects stellar radius in Section V, and then analyze how the TAMS radius regulates supergiant expansion in Section VI.
@@
V Metallicity Effect on Stellar Radius
@@
This section addresses this question: How does metallicity affect the TAMS radius? We begin by analyzing evolutionary tracks, and then apply homology relations to perform a dimensional analysis.
V.1 Metallicity Effect on Stellar Evolutionary Tracks
We have shown that $\eta_{\rm CNO}$ and $\zeta_{\kappa}$—both metallicity-dependent factors—influence whether a star evolves into an RSG by modifying its $R_{\rm TAMS}$. To examine how each parameter affects $R_{\rm TAMS}$, we present the evolutionary tracks for two subsets of the Grid (c) models: one in which $\eta_{\rm CNO}$ varies while $\zeta_{\kappa}$ is fixed (Figure 7), and another in which $\zeta_{\kappa}$ varies while $\eta_{\rm CNO}$ is held constant (Figure 8).
@@ V.1 Metallicity Effect on Stellar Evolutionary Tracks
In Figure 7, the ZAMS positions are not significantly affected by changes in $\eta_{\rm CNO}$, but the subsequent evolutionary tracks exhibit a systematic shift in $T_{\rm eff}$ across different values of $\eta_{\rm CNO}$. Throughout the main-sequence phase and up to the TAMS, higher values of $\eta_{\rm CNO}$ result in lower $T_{\rm eff}$, while the luminosities remain similar. Consequently, stars with enhanced H-burning rates attain larger $R_{\rm TAMS}$.
In Figure 8, variations in $\zeta_{\kappa}$ produce noticeable differences in luminosity from the ZAMS onward, accompanied by moderate shifts in $T_{\rm eff}$. In general, a higher opacity (larger $\zeta_{\kappa}$) leads to a lower luminosity. The ZAMS locations lie nearly parallel to iso-radius lines, indicating that the stellar radius is only weakly sensitive to opacity at this stage. As models evolve, the higher opacity gradually decreases $T_{\rm eff}$. By the TAMS—just before expansion into the supergiant phase—models with higher opacity exhibit systematically lower $T_{\rm eff}$ and larger radii.
@@
These results show that during the main-sequence phase, both increased opacity and a higher H-burning rate shift stellar evolutionary tracks toward cooler effective temperatures and larger radii on the HR diagram. At higher metallicities—where both opacity and the H-burning rate are enhanced—stars are therefore expected to develop systematically larger $R_{\rm TAMS}$. This cumulative metallicity effect is illustrated in Figure 9, which presents evolutionary tracks of $25\,M_{\odot}$ models with different metallicities from Grid (a). Although the ZAMS positions display a relatively irregular distribution with $Z$, the tracks evolve towards a clearer metallicity trend near the TAMS: higher-$Z$ models reach lower $T_{\rm eff}$ at the TAMS, producing systematically larger $R_{\rm TAMS}$. This behavior reflects the influence of metallicity through its combined effects on opacity and nuclear burning.
@@
Figure 7: Evolutionary tracks of $25\,M_{\odot}$, $Z=0.001$ stars with different $\eta_{\rm CNO}$ (subset of Grid (c)) on the HR diagram. Dashed lines indicate constant-radius contours at 10, 100, and $1000\,R_{\odot}$. Black triangles mark the ZAMS, black circles indicate the TAMS, and black stars denote the maximum stellar radius reached during the core He-burning phase. Tracks that evolve into RSGs are shown in red, while those that remain BSGs are shown in blue.
@@
Figure 8: Evolutionary tracks of $25\,M_{\odot}$, $Z=0.001$ stars with different $\zeta_{\kappa}$ (subset of Grid (c)) on the HR diagram. Dashed lines indicate constant-radius contours at 10, 100, and $1000\,R_{\odot}$. Black triangles mark the ZAMS, black circles indicate the TAMS, and black stars denote the maximum stellar radius reached during the core He-burning phase. Tracks that evolve into RSGs are shown in red, while those that remain BSGs are shown in blue.
Figure 9: Evolutionary tracks of $25\,M_{\odot}$ stars with different metallicities (subset of Grid (a)) on the HR diagram. Dashed lines indicate constant-radius contours at 10, 100, and $1000\,R_{\odot}$. Black triangles mark the ZAMS, black circles indicate the TAMS, and black stars denote the maximum stellar radius reached during the core He-burning phase. Tracks that evolve into RSGs are shown in red, while those that remain BSGs are shown in blue.
V.2 Homology Relation between Metallicity and Radius